OpenStudy (anonymous):

help plzzzz Perform the indicated operation. Be sure the answer is reduced. 4k^2/(2k+3)x(2k^2-5k-12)/(8k-32) k/2 k^2/2 2k^2

OpenStudy (anonymous):

\[\frac{ 4k }{ 2k+3 }\times \frac{ 2k ^{2}-5k-12 }{ 8k-32 }\] \[\frac{ k }{ 2 }\]

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

you just need to factorise the numeration and denominator on the right and then lots cancels